# PRobELM: Plausibility Ranking Evaluation for Language Models

### v1

@@ Appendix D More details on PRobELM
Question: What position does <subject>hold? Answer: <object>.
P1532
country for sport
@@
Question: Who represents <object>in sports competitions? Answer: <subject>.
P451
unmarried partner
Question: Who is the unmarried partner of <subject>? Answer: <object>.
P54
member of sports team
@@
Question: <subject>is a member of which sports team? Answer: <object>.
P800
notable work
Question: What is a notable work of <subject>? Answer: <object>.
P551
residence
Question: Where is the residence of <subject>? Answer: <object>.
P131
located in the administrative territorial entity
@@
Question: In which administrative territorial entity is <subject>located? Answer: <object>.
P106
occupation
Question: What is the occupation of <subject>? Answer: <object>.
P69
educated at
Question: Where was <subject>educated? Answer: <object>.
P509
cause of death
Question: What was the cause of death of <subject>? Answer: <object>.
P102
member of political party
@@
Question: Which political party is <subject>a member of? Answer: <object>.
P19
place of birth
Question: Where is the place of birth of <subject>? Answer: <object>.
P115
home venue
Question: What is the home venue of <subject>? Answer: <object>.
P1001
applies to jurisdiction
@@
Question: To which jurisdiction does <object>apply? Answer: <subject>.
P840
narrative location
Question: What is the narrative location of <subject>? Answer: <object>.
P108
employer
Question: Who is the employer of <subject>? Answer: <object>.
P57
director
Question: Who is the director of <object>? Answer: <subject>.
P2416
sports discipline competed in
@@
Question: In which sports discipline does <subject>compete? Answer: <object>.
P400
platform
Question: On which platform is <subject>available? Answer: <object>.
P1433
published in
Question: Where was <subject>published? Answer: <object>.
P1056
product or material produced
@@
Question: What product or material does <subject>produce? Answer: <object>.
P9071
character type
@@
Question: What type of character is <subject>characterized as? Answer: <object>.
P4100
parliamentary group
@@
Question: Which parliamentary group is <subject>a member of? Answer: <object>.
P937
work location
Question: What is the work location of <subject>? Answer: <object>.
P1066
student of
Question: Who is <subject>a student of? Answer: <object>.
P1535
used by
Question: Who uses <object>? Answer: <subject>.
P6
head of government
Question: Who is the head of government of <object>? Answer: <subject>.
P2283
use
Question: What is <subject>used for? Answer: <object>.
P812
academic major
Question: What is <subject>’s academic major?Answer: <object>.
P1416
affiliation
@@
Question: What is <subject>affiliated with? Answer: <object>.
P2522
victory
Question: In which <object>did <subject>achieve a victory? Answer: <object>.
P607
conflict
Question: In which conflict was <subject>involved? Answer: <object>.
P749
parent organization
@@
Question: What parent organization is <subject>a part of? Answer: <object>.
P2283
uses
Question: What does <subject>use? Answer: <object>.
P802
student
Question: At which institution is <subject>a student? Answer: <object>.
P119
place of burial
@@
Question: Where is <subject>buried? Answer: <object>.
P2842
place of marriage
@@
Question: Where was <subject>married? Answer: <object>.
P286
head coach
Question: Who is the head coach of <object>? Answer: <subject>.
P2541
operating area
Question: What is the operating area of <subject>? Answer: <object>.
P1441
present in work
Question: In which work is <subject>present? Answer: <object>.
P2650
interested in
Question: What is <subject>interested in? Answer: <object>.
P1027
conferred by
Question: Who conferred <object>? Answer: <subject>.
P3300
musical conductor
Question: Who is the musical conductor of <object>? Answer: <subject>.
P2715
elected in
Question: In what was <subject>elected? Answer: <object>.
P2937
parliamentary term
@@
Question: During which parliamentary term did <subject>serve? Answer: <object>.
P1399
convicted of
Question: What was <subject>convicted of? Answer: <object>.
P1686
for work
@@
Question: What is <subject>used for in terms of work? Answer: <object>.
P1196
manner of death
Question: What was the manner of death of <subject>? Answer: <object>.
P2632
place of detention
@@
Question: Where was <subject>detained? Answer: <object>.
P991
successful candidate
@@
Question: In which election or selection was <subject>the successful candidate? Answer: <object>.
P2443
stage reached
@@
Question: What stage did <subject>reach? Answer: <object>.
P6872
has written for
Question: For whom or what has <subject>written? Answer: <object>.

### v2

@@ Appendix D More details on PRobELM
P39 position held Question: What position does <subject>hold? Answer: <object>.
@@
P1532 country for sport Question: Who represents <object>in sports competitions? Answer: <subject>.
P451 unmarried partner Question: Who is the unmarried partner of <subject>? Answer: <object>.
@@
P54 member of sports team Question: <subject>is a member of which sports team? Answer: <object>.
P800 notable work Question: What is a notable work of <subject>? Answer: <object>.
P551 residence Question: Where is the residence of <subject>? Answer: <object>.
@@
P131 located in the administrative territorial entity Question: In which administrative territorial entity is <subject>located? Answer: <object>.
P106 occupation Question: What is the occupation of <subject>? Answer: <object>.
P69 educated at Question: Where was <subject>educated? Answer: <object>.
P509 cause of death Question: What was the cause of death of <subject>? Answer: <object>.
@@
P102 member of political party Question: Which political party is <subject>a member of? Answer: <object>.
P19 place of birth Question: Where is the place of birth of <subject>? Answer: <object>.
P115 home venue Question: What is the home venue of <subject>? Answer: <object>.
@@
P1001 applies to jurisdiction Question: To which jurisdiction does <object>apply? Answer: <subject>.
P840 narrative location Question: What is the narrative location of <subject>? Answer: <object>.
P108 employer Question: Who is the employer of <subject>? Answer: <object>.
P57 director Question: Who is the director of <object>? Answer: <subject>.
@@
P2416 sports discipline competed in Question: In which sports discipline does <subject>compete? Answer: <object>.
P400 platform Question: On which platform is <subject>available? Answer: <object>.
P1433 published in Question: Where was <subject>published? Answer: <object>.
@@
P1056 product or material produced Question: What product or material does <subject>produce? Answer: <object>.
@@
P9071 character type Question: What type of character is <subject>characterized as? Answer: <object>.
@@
P4100 parliamentary group Question: Which parliamentary group is <subject>a member of? Answer: <object>.
P937 work location Question: What is the work location of <subject>? Answer: <object>.
P1066 student of Question: Who is <subject>a student of? Answer: <object>.
P1535 used by Question: Who uses <object>? Answer: <subject>.
P6 head of government Question: Who is the head of government of <object>? Answer: <subject>.
P2283 use Question: What is <subject>used for? Answer: <object>.
P812 academic major Question: What is <subject>’s academic major?Answer: <object>.
@@
P1416 affiliation Question: What is <subject>affiliated with? Answer: <object>.
P2522 victory Question: In which <object>did <subject>achieve a victory? Answer: <object>.
P607 conflict Question: In which conflict was <subject>involved? Answer: <object>.
@@
P749 parent organization Question: What parent organization is <subject>a part of? Answer: <object>.
P2283 uses Question: What does <subject>use? Answer: <object>.
P802 student Question: At which institution is <subject>a student? Answer: <object>.
@@
P119 place of burial Question: Where is <subject>buried? Answer: <object>.
@@
P2842 place of marriage Question: Where was <subject>married? Answer: <object>.
P286 head coach Question: Who is the head coach of <object>? Answer: <subject>.
P2541 operating area Question: What is the operating area of <subject>? Answer: <object>.
P1441 present in work Question: In which work is <subject>present? Answer: <object>.
P2650 interested in Question: What is <subject>interested in? Answer: <object>.
P1027 conferred by Question: Who conferred <object>? Answer: <subject>.
P3300 musical conductor Question: Who is the musical conductor of <object>? Answer: <subject>.
P2715 elected in Question: In what was <subject>elected? Answer: <object>.
@@
P2937 parliamentary term Question: During which parliamentary term did <subject>serve? Answer: <object>.
P1399 convicted of Question: What was <subject>convicted of? Answer: <object>.
@@
P1686 for work Question: What is <subject>used for in terms of work? Answer: <object>.
P1196 manner of death Question: What was the manner of death of <subject>? Answer: <object>.
@@
P2632 place of detention Question: Where was <subject>detained? Answer: <object>.
@@
P991 successful candidate Question: In which election or selection was <subject>the successful candidate? Answer: <object>.
@@
P2443 stage reached Question: What stage did <subject>reach? Answer: <object>.
P6872 has written for Question: For whom or what has <subject>written? Answer: <object>.